OpenStudy (anonymous):

Give the slope and y-intercept for the equation below. 10x + 3y = 1 m =? y-intercept =?

OpenStudy (anonymous):

m= -10/3 y-int = 1/3

OpenStudy (aroub):

3y=-10x+1 y=-10/3x+1/3 slope=-10/3 y-intercept=1/3

OpenStudy (anonymous):

aroub you are freakin amazing! thank you.

OpenStudy (aroub):

lol! no prob! anytime!:D